Please show me how to work these problems:

1. The following table shows different bundles that are located in the same indifference curve. Bundle x2 x1 A 161 4 B 114 18 C 79 35 D 65 74 E 22 110 What is the MRS from A to E? (beware of signs)

2. If P2=38, P1=3 and I=15,826 and a tax=52% on x2. What would be the slope of the budget line? (assume the tax on x2 is implemented)
Answer #1

1. MRS from A to E= -1.3

Reason- Point A (x1,x2)=4,161

Point B (x1,x2)= 110,22

MRS= ∆x2/∆x1= (22-161)/(110-4)=-139/106= -1.3

2. Slope of Budget line=-0.05

Reason- New P2=(1+0.52)*38= 57.76

P1=3

Slope=-P1/P2= -3/57.76= -0.05